OpenStudy (anonymous):

what's the inverse of f(x)=12x+11

OpenStudy (anonymous):

y=12x+11 show this function as x=...

OpenStudy (anonymous):

so x-11 over 12?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=f%28x%29%3D12x%2B11 got the answer with the slope

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so easy: first put f(x)=y second : y=12x+11 then you need to make x alone in one side : y-11=12x ==> (y-11)/12=x at the end you must change the place of x to y y=(x-11)/12 remember in f(x) we gave number to x from domain but in f(x)inverse we gave number to x from range of f(x)

OpenStudy (anonymous):

f inverse(x) = (x-11)/12